OpenStudy (anonymous):

if the sum of interior angles of a polygon is 540 degrees, how many sides does the polygon have a. 5 b. 6 c. 7 d. 8

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

awh, good question. remember this equation: (n-2)*180 = sum of interrior angles? this time we know that the sum = 540, so we need to solve for 'n' this time

OpenStudy (anonymous):

so subtract 180 on both sides, or divide? thats where im stuck

OpenStudy (anonymous):

i divided and added two to both sides and got 5. is that correct?

OpenStudy (anonymous):

"so subtract 180 on both sides, or divide? thats where im stuck" the 180 is times, so we have to divide both sides by 180, great question ^_^

OpenStudy (anonymous):

(n-2)*180 = 540 (n-2)*180/180 = 540/180 (n-2) = 540/180 n-2 = 540/180 n-2 +2 = (540/180) +2 n = (540/180) +2

OpenStudy (anonymous):

5 is what i got. and BTW thank you so much you've been a huge help because you actually are helping me understand, not just giving me answers. i appreciate that

OpenStudy (anonymous):

yep, 5 is right, sorry, i didn' see your post that said 5 before ^_^ youre welcome! u got anymore question?

OpenStudy (anonymous):

yeah a few i will just tag you in them :)